A student throws a water balloon vertically downward from the top of a building. The balloon leaves the
thrower's hand with a speed of 6.00 m/s. Air resistance may be ignored, so the water balloon is in free
fall after it leaves the thrower's hand.
1.What is its speed after falling for 2.00 s?
2.How far does it fall in 2.00 s?
1
Expert's answer
2019-10-04T12:55:41-0400

1.


"v=v_0+gt=6+9.8(2)=25.6\frac{m}{s}"

2.


"h=v_0t+0.5gt^2=6(2)+0.5(9.8)(2)^2=31.6\ m"